# MOKA: Open-Vocabulary Robotic Manipulation through Mark-Based Visual Prompting

### v1

@@ V Experiments
The primary goal of our experiments is to validate and analyze the behavior of MOKA on open-vocabulary tasks with a wide range of objects and skills. We design our experiments to investigate the following questions:
  • Can MOKA  effectively perform affordance and motion reasoning on 2D images for open-vocabulary tasks?
@@
Each of our evaluation tasks has two stages, and each task involves a variety of object interaction and tool-use scenarios. A summary of our evaluation tasks can be found in Tab. II. We also tested additional open-vocabulary tasks to demonstrate our method, which can be found in Appendix C.
@@ Methods
TABLE I: Success rate of our method and baselines. Across 4 tasks, each consists of 2 subtasks, MOKA consistently achieves superior performances. We also demonstrated that the performance can be further bootstrapped through physical interactions in the environment using policy distillation and in-context learning.
@@
Figure 4: Failure analysis. We breakdown the failure cases of the variants of our approach using zero-shot prediction, distilled policies, and in-context learning. We breakdown the failures into reasoning failures (caused by errors in the affordance prediction) and execution failures (caused by low-level motions).
@@ V-A Experimental Setup
We compare MOKA with Code-as-Policies [25] and VoxPoser [17], two baselines that also enable zero-shot execution of open-vocabulary tasks:. Code-as-Policies provides a framework for language model-generated programs executed on robotic systems by prompting with code examples. For a fair comparison, we provide the two baselines with the task description in the code comments, with an additional 40 lines of code prompts providing example usage, as in the original implementation [25]. Similarly, VoxPoser [17] also provides code examples to large language models to build a 3D voxel map of value functions. For VoxPoser, we reuse the example prompts and planning pipeline in the original implementation, and use the same hyper-parameters to create the voxel value map. For both baselines, we only adapt the perception modules for fair comparisons, while retaining the functionality of the other components.
We evaluate MOKA in both zero-shot and in-context learning settings, and refer to them as MOKA zero-shot and MOKA in-context, respectively. For the in-context learning setting, we provide two examples to GPT-4V with instructions and annotated images before querying information about the current task. The annotated images are collected in scenes with object and instruction variations. Some in-context prompt examples can be found in Appendix.
The successful trajectories generated by MOKA can also serve as demonstration data for other learning-based methods. By simply training a model using the successful trajectories generated from MOKA, we can also “distill” the knowledge from a VLM to a learned policy. To achieve this goal, we employ a recent open-sourced robot foundation model Octo [35], which is pre-trained on a diverse mix of 800K robot trajectories [36] and can be easily fine-tuned to new tasks. Octo [35] supports language instructions or goal images as task specifications, and generates actions observations with a transformer-based diffusion policy architecture. It also provides code examples to easily perform fine-tuning. To construct our fine-tuning dataset, we collect 50 successful trajectories for each task with language annotations. We then adopt its latest base model checkpoint [^1] https://huggingface.co/rail-berkeley/octo-base, and finetune the full model with recipe provided in the official instruction [^2] https://github.com/octo-models/octo. The performance of fine-tuned can be found in Sec. V-B with more implementation details in Appendix B-E. Although MOKA is a training-free method, it can be leveraged for data-driven approaches to provide training data of open-vocabulary tasks. We refer to this approach as MOKA-Distilled.
@@ V-B Quantitative Evaluation
Our quantitative evaluation results across 4 tasks are illustrated in Tab. I. For each task, we report the number of successes out of 10 trials. As shown in the the table, MOKA achieves state-of-the-art performance at each subtask of the 4 tasks (totally 8 subtasks), with consistent improvements using in-context learning. On most of the tasks, VoxPoser [17] has similar performance with MOKA (zero-shot), except for subtask 2 of table wiping (which is a tool-use task). Additionally, the task success rates can be sensitive to the resolution of the voxel map, which requires some hyperparameter tuning. Unlike the baselines, MOKA can work well without example prompts, or can achieve better performance with just two clean and intuitive example prompts.
We also observed that we can obtain an effective end-to-end policy by finetuning a pre-trained robot policy [35] using the successful trajectories generated by MOKA. This suggests that the generated data is of high quality, and thus can be used as demonstration data for open-vocabulary tasks. The fact that distilling the successful MOKA trials actually improves the overall performance of the system suggests the feasibilty of using MOKA as a “data generator”, where the zero-shot MOKA method is used to bootstrap a continuous improvement system. This is an exciting direction to explore in future work.
Failure breakdown We analyze the failure cases of MOKA. Trajectories with wrong predictions from the GPT-4V are counted as reasoning failures. The following failure cases are illustrated in the top row of Fig. 4, including grasping the broom upside down due to confusing the the grasp point with the function point, pointing the room to the wrong direction due to wrong target angle, pressing the hinge of the laptop due to the wrong target point. The trajectories with desired VLM prediction but fail to execute successfully are counted as execution failures. The following failure cases are illustrated in the bottom row of Fig. 4. From left to right, the failures include the gripper narrowly misses the button, the filler getting disassembled in the middle of the trajectory, and the cable slipped through the gripper fingers.
As shown in the figure, both policy distillation and in-context learning reduces the total failures. Using the distilled policy, the VLM is no longer part of the pipeline. Therefore there is no reasonoing failures in this case. But the execution failures increases since the policy does not realy on the open-vocabulary detection and segmentation.
@@ V-C Qualitative Evaluation
In Fig. 3, we provide six examples of the annotated marks, generated motions, and the unrolled trajectories of MOKA. In each of these examples, based on the annotated marks, MOKA successfully selected the keypoints and waypoints for generating the desired motions. These examples demonstrate that our approach can be applied to various skills including pressing, closing, rearranging, tool use, etc.
Figure 5: Robustness analysis. We analyze MOKA’s robustness with respect to various instructions, initial arrangements, and objects. Each column in the image uses the same language instruction and similar initial arrangements of objects. The two rows involve different objects.
We analyze the robustness of MOKA with respect to instruction variations, initialization variations, and object variations. Fig. 5 provides a pictorial illustration of MOKA’s predictions under different instructions and observations. Each image is queried with the language prompt on the top within the same column. The examples in the same column share similar initialization object positions and orientations. The example in the first row uses the same set of objects while the second row involves objects of alternative geometries, colors, and materials. Some of these objects are also deformable or transparent. Fig. 5 demonstrates consistent point predictions within rows and columns, indicating that MOKA is robust to the changes of the language instructions, initialization, and objects for the same task.
@@ VI Conclusion and Discussion
In this paper, we proposed MOKA, a simple and effective visual prompting method that leverages VLMs for robot manipulation. By representing manipulation tasks with point-based affordances, we convert the motion generation process to a visual question-answering problem that VLMs can solve. MOKA provides a general and flexible framework that can intuitively and effectively harness VLM to generate point-based motion for a wide range of open-vocabulary tasks, while preserving the visual reasoning capabilities of VLMs. Our experiments demonstrate the effectiveness and robustness of MOKA across multiple tasks in both zero-shot and in-context learning manners. As far as we know, MOKA is the very first method that leverages visual prompting on VLMs for open-vocabulary robot manipulation.
The diversity and difficulty of solvable tasks for MOKA depend on the capabilities of VLMs. For example, we mostly evaluate MOKA in 4-DoF table-top manipulation tasks, as current VLMs are not capable of reliably predicting 6-DoF motions. It will also be hard to extend MOKA to more challenging scenarios, e.g. bi-manual manipulation tasks, where the coordination of two arms requires more complex 3D understanding. In addition, the dependency of querying third-party VLM APIs also constraints us from applying MOKA to dynamic tasks, with low latency tolerance. We believe the above limitations can be conquered with more capable and accessible VLMs in the future.
In our experiments, we also demonstrate that the trajectories generated by VLM can be further used to train a learning-based policy that yields better performance. We hope MOKA can inspire future research toward designing more efficient and scalable algorithms to collect and bootstrap on real-world open-vocabulary datasets more autonomously.
VII Acknowledgement
Fangchen Liu is supported in part by NSF under the NSF AI4OPT Center. We would like to thank Karl Pertsch, Oleg Rybkin, and Zheyuan Hu for providing feedback on early drafts of the manuscript. We also want to thank Sudeep Dasari, Qiyang Li, Yide Shentu, and Homer Walke for valuable suggestions on the method and task design.
@@ A-A Environment
In addition, we also set up the front camera and the wrist camera for more complete observation to distill the collected robot experiences to the learned Octo policy [35]. The action space of the robot is 7-dimensional, consisting of the 6DoF end-effector twist defined in the Cartesian space with an additional dimension of gripper position. Each task can consist of multiple stages, each terminates within a fix horizon of 100 steps.
@@ A-B Task Design
We design various table-top manipulation tasks with a diverse set of daily objects. Tab. III shows the language description of the full list of tasks.
@@ 2. Insert the roses into the vase
We provide the comparative evaluation on the top four tasks (Table Wiping, Laptop Packing, Gift Preparation and Untrasound Cleaning) in the main paper, with additional results of MOKA on other 4 tasks in our supplementary video.
@@ B-A High-level Task Reasoning
Given the initial observation image $s_{0}$ and the language task description $l$, we first query the VLM $\mathcal{M}$ with the language instruction to produce the decomposed subtask, which contains the structured information for the $K$ subtasks, including descriptions of objects and desired motion. We provide the high-level reasoning we used across all the tasks in Tab. IV.
The input request contains: A string describes the multi-stage task. An image of the current table-top environment captured from a top-down camera. List of object candidates in the scene. The output response is a list of dictionaries in the JSON form. Each dictionary specifies the information of a subtask, following the correct order of executing the subtasks to solve the input task. Each dictionary contains these fields: instruction: A string to describe the subtask in natural language forms. object_normal-_\__grasped: A string to describe the name of the object that the robot gripper will hold in hand while executing the task (e.g., the object to be picked or used as a tool to interact with other objects). This field can be empty if there is no such object involved in this subtask. object_normal-_\__unattached: A string to describe the name of the object that the robot gripper will interact with directly or via another object without holding it in hand (e.g., the object to be touched by the tool, the target object where “object__\__grasped” will be moved onto). This field can be empty if there is no such object involved in this subtask. motion_normal-_\__direction: A string to describe the direction of the robot gripper motion while performing the task (e.g., “from right to left”, “backward”, “downward”).
@@ B-B Point-based Affordance Proposal
Keypoint proposal. We leverage GroundedSAM [43] to extract segmentation masks conditioned on a text prompt, which is designed to be a string of object names involved in the current subtask (e.g. “trash, broom”). Given an image of current observation and such a text prompt about the involved object names, we first employ GroundingDINO [27] to generate bounding boxes for the objects by conditioning on the text prompt. Later, the annotated boxes given by GroundingDINO [27] serve as the box prompts for SAM [18] to generate corresponding segmentation masks for the objects. We define the keypoints of each object to be a set of boundary keypoints with a center keypoint. To extract them, we sample $K$ points on the contour of each object using farthest point sampling [37], and also include the geometric mean point of the object segmentation mask. We then plot these $K+1$ keypoints on the 2D image as our keypoint proposals to the VLM.
@@ B-C Low-level Motion Reasoning
Please describe the robot gripper’s motion to solve the task by selecting keypoints and waypoints. The input request contains: The task information with these fields: instruction: The task in natural language forms. object_normal-_\__grasped: The object that the robot gripper will hold in hand while executing the task. object_normal-_\__unattached: The object that the robot gripper will interact with either directly or via another object without holding it in hand. motion_normal-_\__direction: The motion direction of the robot gripper or the in-hand object while performing the task. An image of the current table-top environment captured from a top-down camera, annotated with a set of visual marks: candidate keypoints on object_normal-_\__grasped: Red dots marked as $P[i]$ on the image. candidate keypoints on object_normal-_\__unattached: Blue dots marked as $Q[i]$ on the image. grid for waypoints: Grid lines that uniformly divide the images into tiles. The grid equally divides the image into columns marked as $a,b,c,d,e$ from left to right and rows marked as 1, 2, 3, 4, 5 from bottom to top.
@@
The motion consists of an optional grasping phase and a manipulation phase, specified by grasp_normal-_\__keypoint, function_normal-_\__keypoint, target_normal-_\__keypoint, pre_normal-_\__contact_normal-_\__waypoint, and post_normal-_\__contact_normal-_\__waypoint. The definitions of these points are: grasp_normal-_\__keypoint: The point on “object__\__grasped” indicates the part where the robot gripper should hold. function_normal-_\__keypoint: The point on “object__\__grasped” indicates the part that will make contact with “object__\__unattached”. target_normal-_\__keypoint: If the task is pick-and-place, this is the location where “object__\__grasped” will be moved to. Otherwise, this is the point on “object__\__unattached” indicating the part that will be contacted by “function__\__keypoint”. pre_normal-_\__contact_normal-_\__waypoint: The waypoint in the free space that the functional point moves to before making contact with the “target__\__keypoint”. post_normal-_\__contact_normal-_\__waypoint: The waypoint in the free space that the functional point moves to after making contact with the “target__\__keypoint”.
@@
After that, we can specify the output format and further explain the role of each field using prompt VII. We can also provide some step-by-step guidance about how the reasoning procedure should be done (similar to chain-of-thought prompting [50]). Our complete low-level reasoning prompts are covered in Tab. V,  VI and  VII.
The response should be a dictionary in JSON form, which contains: grasp_normal-_\__keypoint: Selected from candidate keypoints marked as $P[i]$ on the image. This will be empty if and only if object__\__grasped is empty. function_normal-_\__keypoint: Selected from candidate keypoints marked as $P[i]$ on the image. This will be empty if and only if object__\__grasped or object__\__unattached is empty. target_normal-_\__keypoint: Selected from keypoint candidates marked as $Q[i]$ on the image. This will be empty if and only if object__\__unattached is empty. pre_normal-_\__contact_normal-_\__tile: The tile that the pre-contact waypoint should be in. This is selected from candidate tiles marked on the image. post_normal-_\__contact_normal-_\__tile: The tile that post-contact waypoint should be in. This is selected from candidate tiles marked on the image. pre_normal-_\__contact_normal-_\__height: The height of pre-contact waypoint as one of the two options “same” or “above” (same or higher than the height of making contact with target keypoint). post_normal-_\__contact_normal-_\__height: The height of post-contact waypoints as one of the two options “same” or “above”. target_normal-_\__angle: Describe how the object should be oriented during this motion in terms of the axis pointing from the grasping point to the function point. Think about this problem step by step and explain the reasoning steps. First, choose grasp__\__keypoint, function__\__keypoint, and target__\__keypoint on the correct parts of the objects. Next, describe which tile the target__\__keypoint is located in. Then choose pre__\__contact__\__tile, post__\__contact__\__tile, pre__\__contact__\__height, post__\__contact__\__height such that the resultant motion from pre-contact waypoint to target keypoint, then to post-contact waypoint in 3D follows the ”motion__\__direction” input. Remember that the columns are marked as ’a’, ’b’, ’c’, ’d’, ’e’ from left to right, and the rows are marked as 1, 2, 3, 4, 5 from bottom to top.

### v3

@@ V Experiments
The primary goal of our experiments is to validate and analyze the behavior of MOKA on open-world tasks with a wide range of objects and skills. We design our experiments to investigate the following questions:
  • Can MOKA  effectively reason about affordances based on the images for unseen tasks?
@@
Each of our evaluation tasks has two stages, and each task involves a variety of object interaction and tool-use scenarios. A summary of our evaluation tasks can be found in TABLE II. We also tested additional open-world tasks to demonstrate our method, which can be found in Appendix C.
@@ Methods
TABLE I: Success rate of our method and baselines. Across four tasks, each consists of 2 subtasks, MOKA consistently achieves superior performances. We also demonstrated that the performance can be further bootstrapped through physical interactions in the environment using policy distillation and in-context learning.
@@
Figure 4: Failure analysis. We break down the failure cases of the variants of our approach using zero-shot prediction, distilled policies, and in-context learning. We break down the failures into reasoning failures (caused by errors in the affordance prediction) and execution failures (caused by low-level motions). We demonstrate how policy distillation and in-context learning reduce failure cases.
@@ V-A Experimental Setup
We compare MOKA with Code-as-Policies [34] and VoxPoser [23], two baselines that also enable zero-shot execution of open-world tasks:. Code-as-Policies provides a framework for language model-generated programs executed on robotic systems by prompting with code examples. For a fair comparison, we provide the two baselines with the task description in the code comments, with an additional 40 lines of code prompts providing example usage, as in the original implementation [34]. Similarly, VoxPoser [23] also provides code examples to large language models to build a 3D voxel map of value functions. For VoxPoser, we reuse the example prompts and planning pipeline in the original implementation and use the same hyper-parameters to create the voxel value map. For both baselines, we only adapt the perception modules for fair comparisons while retaining the functionality of the other components.
We evaluate MOKA in both zero-shot and in-context learning settings and refer to them as MOKA zero-shot and MOKA in-context, respectively. For the in-context learning setting, we provide two examples to GPT-4V with instructions and annotated images before querying information about the current task. The annotated images are collected in scenes with object and instruction variations. Some in-context prompt examples can be found in the Appendix.
The successful trajectories generated by MOKA can also serve as demonstration data for other learning-based methods. By simply training a model using the successful trajectories generated from MOKA, we can also “distill” the knowledge from a VLM to a learned policy. To achieve this goal, we employ a recent open-sourced robot foundation model Octo [46], which is pre-trained on a diverse mix of 800K robot trajectories [47] and can be easily fine-tuned to new tasks. Octo [46] supports language instructions or goal images as task specifications and takes current observations to generate actions with a transformer-based diffusion policy architecture. It also provides code examples to easily perform fine-tuning. To construct our fine-tuning dataset, we collect 50 successful trajectories for each task with language annotations. We then adopt its latest base model checkpoint [^1] https://huggingface.co/rail-berkeley/octo-base and finetune the full model with the recipe provided in the official instruction [^2] https://github.com/octo-models/octo. The performance of fine-tuned can be found in Sec. V-B with more implementation details in Appendix B-E. Although MOKA is a training-free method, it can be leveraged for data-driven approaches to provide training data for open-world tasks. We refer to this approach as MOKA-Distilled.
@@ V-B Quantitative Evaluation
Our quantitative evaluation results across 4 tasks are illustrated in TABLE I. For each task, we report the number of successes out of 10 trials. As shown in the table, MOKA achieves state-of-the-art performance at each subtask of the 4 tasks (total 8 subtasks), with consistent improvements using in-context learning. On most of the tasks, VoxPoser [23] has similar performance with MOKA (zero-shot), except for subtask 2 of table wiping (which is a tool-use task). Additionally, the task success rates can be sensitive to the resolution of the voxel map, which requires some hyperparameter tuning. Unlike the baselines, MOKA can work well without example prompts or can achieve better performance with just two clean and intuitive example prompts.
We also observed that we can obtain an effective end-to-end policy by fine-tuning a pre-trained robot policy [46] using the successful trajectories generated by MOKA. This suggests that the generated data is of high quality and thus can be used as demonstration data for open-world tasks. The fact that distilling the successful MOKA trials actually improves the overall performance of the system suggests the feasibility of using MOKA as a “data generator”, where the zero-shot MOKA method is used to bootstrap a continuous improvement system. This is an exciting direction to explore in future work.
Failure breakdown We analyze the failure cases of MOKA. Trajectories with wrong predictions from the GPT-4V are counted as reasoning failures. The following failure cases are illustrated in the top row of Fig. 4, including grasping the broom upside down due to confusing the grasp point with the function point, pointing the room in the wrong direction due to the wrong target angle, pressing the hinge of the laptop due to the wrong target point. The trajectories with desired VLM prediction but fail to execute successfully are counted as execution failures. The following failure cases are illustrated in the bottom row of Fig. 4. From left to right, the failures include the gripper narrowly missing the button, the filler getting disassembled in the middle of the trajectory, and the cable slipping through the gripper fingers.
As shown in the figure, both policy distillation and in-context learning reduce the total failures. Using the distilled policy, the VLM is no longer part of the pipeline. Therefore, there are no reasoning failures in this case.
@@ V-C Qualitative Evaluation
In Fig. 3, we provide six examples of the annotated marks, generated motions, and the unrolled trajectories of MOKA. In each of these examples, based on the annotated marks, MOKA successfully selected the keypoints and waypoints for generating the desired motions. These examples demonstrate that our approach can be applied to various skills, including pressing, closing, rearranging, tool use, etc.
Figure 5: Robustness analysis. We analyze MOKA’s robustness with respect to various instructions, initial arrangements, and objects. Each column in the image uses the same language instruction and similar initial arrangements of objects. The two rows involve different objects. MOKA  demonstrates to robustly predicts the keypoints and waypoints across a wide range of variations. More examples of other tasks are presented in the Appendix.
We analyze the robustness of MOKA with respect to instruction variations, initialization variations, and object variations. Fig. 5 provides a pictorial illustration of MOKA’s predictions under different instructions and observations. Each image is queried with the language prompt on the top within the same column. The examples in the same column share similar initialization object positions and orientations. The example in the first row uses the same set of objects, while the second row involves objects of alternative geometries, colors, and materials. Some of these objects are also deformable or transparent. Fig. 5 demonstrates consistent keypoint and waypoint predictions within rows and columns, indicating that MOKA is robust to the changes of the language instructions, objects, and initial arrangements for the same task.
@@ VI Conclusion and Discussion
In this paper, we propose MOKA, a simple and effective approach that leverages pre-trained VLMs for open-world robotic manipulation. By specifying motions with point-based affordance representations, we enable VLMs pre-trained on Internet-scale data to solve manipulation tasks in zero-shot and few-shot manners. Using the proposed mark-based visual prompting technique, MOKA converts affordance reasoning into a series of visual question-answering problems that the pre-trained VLMs can solve. Through experiments on a variety of table-top manipulation tasks, we demonstrate the effectiveness and robustness of MOKA. We also demonstrate that the trajectories generated by MOKA can be further used to bootstrap the performance through in-context learning and policy distillation. As far as we know, MOKA is the very first work that leverages visual prompting on pre-trained VLMs for open-world robot manipulation. We hope MOKA will inspire future research towards leveraging vision-language models for open-world robotic control.
MOKA is subject to limitations mostly due to the capabilities of existing VLMs and the current design of the affordance representations. The state-of-the-art VLMs still lack a profound understanding of 3D space, contact physics, and robotic control, which prevents our approach from generating more fine-grained motions in the $SE(3)$ space. The high latency of querying VLMs also makes it challenging for the robot to perform dynamic manipulation and open-loop control. Moreover, we would like to note that, while we demonstrate that a unified point-based affordance representation can be devised to support diverse tasks, the specific design of the representation in this work does not cover all useful skills that a robot can perform. To apply MOKA to more complex scenarios, e.g., bi-manual manipulation and whole-body control, we would need to extend the set of points and include additional attributes in this affordance representation. We believe these limitations will be addressed as more capable vision-language models and robotic foundation models are introduced in the future.
Acknowledgement
This research was supported by the AI Institute, AFOSR FA9550-22-1-0273, and the NSF under IIS-2246811. Fangchen Liu was supported in part by NSF under the NSF AI4OPT Center. We would like to thank Karl Pertsch, Oleg Rybkin, and Zheyuan Hu for providing feedback on early drafts of the manuscript. We also want to thank Sudeep Dasari, Qiyang Li, Yide Shentu, and Homer Walke for valuable suggestions on the experiments and infrastructure.
@@ A-A Environment
In addition, we set up the front camera and the wrist camera for more complete observation to distill the collected robot experiences to the learned Octo policy [46]. The action space of the robot is 7-dimensional, consisting of the 6-DoF end-effector twist defined in the $SE(3)$ space with an additional dimension of gripper position. Each task can consist of multiple stages, each terminates after 100 steps.
@@ A-B Task Design
We design various table-top manipulation tasks with a diverse set of daily objects. TABLE III shows the language description of the full list of tasks.
@@ 2. Insert the roses into the vase
We provide the comparative evaluation on the top four tasks (Table Wiping, Laptop Packing, Gift Preparation and Untrasound Cleaning) in the main paper, with additional results of MOKA on other four tasks in our supplementary video.
@@ B-A High-Level Task Reasoning
Given the initial observation image $s_{0}$ and the language task description $l$, we first query the VLM $\mathcal{M}$ with the language instruction to produce the decomposed subtask, which contains the structured information for the $K$ subtasks, including descriptions of objects and desired motion. We provide the prompt for high-level reasoning in TABLE IV.
{mdframed}
The input request contains:
  • A string describes the multi-stage task.
  • An image of the current table-top environment captured from a top-down camera.
The output response is a list of dictionaries in the JSON form. Each dictionary specifies the information of a subtask, following the correct order of executing the subtasks to solve the input task. Each dictionary contains these fields:
  • instruction: A string to describe the subtask in natural language forms.
  • object__\__grasped: A string to describe the name of the object that the robot gripper will hold in hand while executing the task (e.g.,, the object to be picked or used as a tool to interact with other objects). This field can be empty if there is no such object involved in this subtask.
  • object__\__unattached: A string to describe the name of the object that the robot gripper will interact with directly or via another object without holding it in hand (e.g.,, the object to be touched by the tool, the target object where “object__\__grasped” will be moved onto). This field can be empty if there is no such object involved in this subtask.
  • motion__\__direction: A string to describe the direction of the robot gripper motion while performing the task (e.g.,, “from right to left”, “backward”, “downward”).
@@ B-B Point-based Affordance Proposal
Keypoint proposal. We leverage GroundedSAM [54] to extract segmentation masks conditioned on a text prompt, which is designed to be a string of object names involved in the current subtask (e.g., “trash, broom”). Given an image of current observation and such a text prompt about the involved object names, we first employ GroundingDINO [36] to generate bounding boxes for the objects by conditioning on the text prompt. Later, the annotated boxes given by GroundingDINO [36] serve as the box prompts for SAM [27] to generate corresponding segmentation masks for the objects. We define the keypoints of each object to be a set of boundary keypoints with a center keypoint. To extract them, we sample $K$ points on the contour of each object using farthest point sampling [48], and also include the geometric mean point of the object segmentation mask. We then plot these $K+1$ keypoints on the 2D image as our keypoint proposals to the VLM.
@@ B-C Low-Level Motion Reasoning
{mdframed}
Please describe the robot gripper’s motion to solve the task by selecting keypoints and waypoints.
The input request contains:
  • The task information with these fields:
    • instruction: The task in natural language forms.
    • object__\__grasped: The object that the robot gripper will hold in hand while executing the task.
    • object__\__unattached: The object that the robot gripper will interact with either directly or via another object without holding it in hand.
    • motion__\__direction: The motion direction of the robot gripper or the in-hand object while performing the task.
  • An image of the current table-top environment captured from a top-down camera, annotated with a set of visual marks:
    • candidate keypoints on object__\__grasped: Red dots marked as $P[i]$ on the image.
    • candidate keypoints on object__\__unattached: Blue dots marked as $Q[i]$ on the image.
    • grid for waypoints: Grid lines that uniformly divide the images into tiles. The grid equally divides the image into columns marked as $a,b,c,d,e$ from left to right and rows marked as 1, 2, 3, 4, 5 from bottom to top.
@@
{mdframed}
The motion consists of an optional grasping phase and a manipulation phase, specified by grasp__\__keypoint, function__\__keypoint, target__\__keypoint, pre__\__contact__\__waypoint, and post__\__contact__\__waypoint.
The definitions of these points are:
  • grasp__\__keypoint: The point on “object__\__grasped” indicates the part where the robot gripper should hold.
  • function__\__keypoint: The point on “object__\__grasped” indicates the part that will make contact with “object__\__unattached”.
  • target__\__keypoint: If the task is pick-and-place, this is the location where “object__\__grasped” will be moved to. Otherwise, this is the point on “object__\__unattached” indicating the part that will be contacted by “function__\__keypoint”.
  • pre__\__contact__\__waypoint: The waypoint in the free space that the functional point moves to before making contact with the “target__\__keypoint”.
  • post__\__contact__\__waypoint: The waypoint in the free space that the functional point moves to after making contact with the “target__\__keypoint”.
@@
After that, we can specify the output format and further explain the role of each field using prompt VII. We can also provide some step-by-step guidance about how the reasoning procedure should be done (similar to chain-of-thought prompting [62]).
{mdframed}
The response should be a dictionary in JSON form, which contains:
  • grasp__\__keypoint: Selected from candidate keypoints marked as $P[i]$ on the image. This will be empty if and only if object__\__grasped is empty.
  • function__\__keypoint: Selected from candidate keypoints marked as $P[i]$ on the image. This will be empty if and only if object__\__grasped or object__\__unattached is empty.
  • target__\__keypoint: Selected from keypoint candidates marked as $Q[i]$ on the image. This will be empty if and only if object__\__unattached is empty.
  • pre__\__contact__\__tile: The tile that the pre-contact waypoint should be in. This is selected from candidate tiles marked on the image.
  • post__\__contact__\__tile: The tile that post-contact waypoint should be in. This is selected from candidate tiles marked on the image.
  • pre__\__contact__\__height: The height of pre-contact waypoint as one of the two options “same” or “above” (same or higher than the height of making contact with target keypoint).
  • post__\__contact__\__height: The height of post-contact waypoints as one of the two options “same” or “above”.
  • target__\__angle: Describe how the object should be oriented during this motion in terms of the axis pointing from the grasping point to the function point.
Think about this problem step by step and explain the reasoning steps. First, choose grasp__\__keypoint, function__\__keypoint, and target__\__keypoint on the correct parts of the objects. Next, describe which tile the target__\__keypoint is located in. Then choose pre__\__contact__\__tile, post__\__contact__\__tile, pre__\__contact__\__height, post__\__contact__\__height such that the resultant motion from pre-contact waypoint to target keypoint, then to post-contact waypoint in 3D follows the ”motion__\__direction” input. Remember that the columns are marked as ’a’, ’b’, ’c’, ’d’, ’e’ from left to right, and the rows are marked as 1, 2, 3, 4, 5 from bottom to top.